RC meeting and Run Plan, Wednesday September 21st

From SBS wiki
Jump to: navigation, search


RC Meeting Information

Run Coordinator: Donald Jones

Location: Counting House Conference Room

Time: 3:00PM

Counting House status

Subsystem status reports

· Hall A Status -- Jesse Butler

· Target status/timeline (training?) -- Arun

· Beam line: monitors, vacuum, correctors... -- Simona

· Website/wiki/shift schedule -- Will Tireman/Don

· Tentative run plan document -- Bogdan

· Shift Checklist --Don

· Alarms

· Analyzer -- Andrew

· DAQ -- Alex

· HCal -- Jiwan

· Spectrometer magnets -- Jack

· GEMs/Tracking -- Holly

· Shower Calorimeter -- Arun

· Beam polarization -- Don


  • The RC report:
    • With ongoing target work, we will likely not lock up until Tuesday. He informed the PD at the 1:30 accelerator scheduling meeting and she changed the tentative date to Sept 27 (Tues). He will continue to keep the PD up to date as the situation evolves.
    • He talked with Yves Roblin and Mike Freeman who confirmed that the three days on the schedule labeled as "Lock up" which are Monday through Wednesday, are for unforeseen failures/difficulties and it is likely that moving the lock up to Tuesday will not affect the scheduled delivery of beam on Thursday.
    • Discussed the schedule. We decided to remove all target operator assignments from the calendar for the first full week and give them only to the target expert team.
    • Requested that experts update their sections in the wiki, in particular, the How-Tos
    • Requested that collaborators look at the shift checklist and suggest changes, additions and things that need to be dropped. All shifts will be canceled until Day Tuesday. Will Tireman will oversee the schedule re-arrangement and contact those whose shifts have been cancelled.
    • Showed pictures of clutter in the Counting House back room and asked that it be cleaned up. He will inquire about getting the large fridge back.
  • Jessie is trying to get the target laser enclosure complete and ready for the scheduled inspection at 1:30PM Thursday Sept 22. Activities in the area will require that there be no one working in the target area Thursday morning. He pointed out that no one is authorized to work with the laser in the CANS system so there will be no working in the area with the laser activated until this issue has been solved. Perhaps it is as simple as the overseer of the laser training updating training status in the database (J.P. Chen?).
  • Jack reported that BigBite power supply tripped at half power twice for different reasons: the first time an internal temperature fault and the second an LCW fault. He needs time to troubleshoot when no one is in the area so he can ramp up the magnet. He will try for an hour after 5PM tonight. If he can ramp to full power, he will report any safety issues to the OSP signatories who will add precautions into the OSP for those working in the area with the field on. Arun and Co. have to work in the area with the field on to do field measurements in the target region. Once the OSP is approved they can do that. Jack also asked if the SBS corrector power supply cables were sufficient for GEn. They are limited to 100A. Bogdan said that GEn will not need to go as high as 100A, so the power supply will be limited to a maximum of 100A and the cables left as is. Bogdan will supply the RC with the current settings for the correctors at the four kinematic settings so that they can be published on the wiki for future RCs and shift crews. Bogdan said the first two points are known and he sent them along, but the highest two points still have simulation work to be done.
  • Arun reported on a list of task that have to be completed for the target to be ready. The most critical involve the laser (which has to be approved by the laser safety folks for operation) and the field measurements with SBS and BB on. The field measurements cannot happen until Jack is done with magnet ramping tests and included concerns in the OSP.
  • Simona reported on the raster and that she needs to connect with Bill Gunning to make sure whether or not he wants to be there when manual adjustments to the function generators are made. RC reported that Yves plans to correct distortion of raster with optics to retain circularity of raster. Simona will be doing BPM and BCM calibrations early in the process of establishing beam in the Hall.
  • Yves has a detailed procedure (link) which he and David Flay developed for establishing beam in the hall and in particular assuring that the target and beam are aligned. Arun and Simona will take a look at this plan to ensure it makes sense to them.
  • Bogdan's detailed run plan is now available on wiki